OpenStudy (anonymous):

Anybody wiling to help with algebra 1 for a medal?

OpenStudy (unklerhaukus):

[insert question]

OpenStudy (anonymous):

Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

heres the chart that goes with the problem

OpenStudy (anonymous):

can you open the chart?

OpenStudy (unklerhaukus):

yeah , but its easier to just upload a screen shot

OpenStudy (anonymous):

oh im sorry i didnt even know you could do that lol. can you help? ive been stuck for weeks.

OpenStudy (unklerhaukus):

so x is days right?

OpenStudy (anonymous):

right

OpenStudy (unklerhaukus):

have you done any of Q1?

OpenStudy (anonymous):

i dont even know how to start it

OpenStudy (unklerhaukus):

well A, B, and C are all different expressions for some different stock functions, you have to make stock function D

OpenStudy (anonymous):

okay

OpenStudy (unklerhaukus):

D must have `the second lowest starting price' so What are the starting prices of A, B and C?

OpenStudy (unklerhaukus):

starting prices are the f(0) values

OpenStudy (anonymous):

a=25 b=22 c=30 ??

OpenStudy (unklerhaukus):

yes, that's right

OpenStudy (unklerhaukus):

what could you choose as the starting price of D such that D has `the second lowest starting price'

OpenStudy (anonymous):

so the second lowest starting price would either be 23 or 24?

OpenStudy (unklerhaukus):

yeah you can choose either of those

OpenStudy (anonymous):

okay so 24

OpenStudy (unklerhaukus):

stock function D must also have ` the second highest rate of growth' what are the rates of growth of A, B, C?

OpenStudy (anonymous):

im not sure how to find that. is that the slope or something?

OpenStudy (unklerhaukus):

rates of growth is a similar idea to the slope, but it not quite the same, because the stock values will grow in a curve (not a straight line)

OpenStudy (anonymous):

so how do i find rate of growth of a. b. c?

OpenStudy (unklerhaukus):

exponential growth (as we have here) fits this formula \[f(x)=A(1+r)^x\] the value of the function is \(x\) [the price of the stock] \(A\) is the initial values (starting price) and the growth rate is \(r\)

OpenStudy (unklerhaukus):

\(r\) is often expressed as a percentage

OpenStudy (anonymous):

oh! so for a=1.08? b=? and c=.04

OpenStudy (anonymous):

or is the 1.08 after they added 1? so it would be .08 for a?

OpenStudy (unklerhaukus):

that's right for A , the growth rate is 0.08 , what is this as a percentage?

OpenStudy (anonymous):

8% and 4% for c

OpenStudy (unklerhaukus):

(and yes 0.04 or 4% for the growth rate of C is correct )

OpenStudy (unklerhaukus):

good

OpenStudy (anonymous):

i dont know how to find b rate of change

OpenStudy (unklerhaukus):

the table says f(0) = 22.00 this is the initial value / starting price f(1) = 21.56 and you know that exponential growth has the form \(f(x) = A(1+r)^x\)

OpenStudy (unklerhaukus):

what do you get when you plug the initial values in for \(A\), and use \(x=1\), can you solve this for x?

OpenStudy (anonymous):

yes hold on

OpenStudy (unklerhaukus):

*solve for \(r\)

OpenStudy (anonymous):

f(x)=22(1+r)^1

OpenStudy (anonymous):

22+r^1?

OpenStudy (unklerhaukus):

that's f(1)

OpenStudy (unklerhaukus):

wait what ,

OpenStudy (anonymous):

i thought f(0) was 22

OpenStudy (unklerhaukus):

\[f(1)=22(1+r)^1=21.56\] solve this for \(r\)

OpenStudy (anonymous):

okay so would i distribute 22 first?

OpenStudy (unklerhaukus):

\[f(0)=22(1+r)^0=22(1)^0=22\]

OpenStudy (unklerhaukus):

\[22(1+r)^1=21.56\] divide both sides by 22

OpenStudy (anonymous):

\[(22+r)^{1}=.98\]

OpenStudy (anonymous):

i meant 1 + r

OpenStudy (unklerhaukus):

ok (1+r)^1 = .98 simplify the left hand side

OpenStudy (anonymous):

well the 1st power of 1 is 1 so would it be \[1+r=.98\]

OpenStudy (unklerhaukus):

good, keep going, what is r?

OpenStudy (anonymous):

r=-.02?

OpenStudy (unklerhaukus):

which makes sense, because as x (the days) increase in the table, the f(x) values [stock prices of B are dropping ]

OpenStudy (anonymous):

right

OpenStudy (unklerhaukus):

so you have you growth rates A: r = 0.08 = 8% B: r = -0.02 = C: r = 0.04 = 4% D: r =

OpenStudy (unklerhaukus):

remember the requirement is that stock D has ` the second highest rate of growth'

OpenStudy (unklerhaukus):

@lyndsaysturgeon

OpenStudy (unklerhaukus):

what are you going to choose for the rate of growth for stock D?

OpenStudy (unklerhaukus):

For part 2. Form an equation for the value of stock D like \(f(x)=A(1+r)^x\) with the parameters \(A\) and \(r\), that you have chosen make a table and plot the points

OpenStudy (anonymous):

okay so what would be the percent of -.02?

OpenStudy (unklerhaukus):

0.02 is 2% so -0.02 is -2%

OpenStudy (anonymous):

okay so the percent for d would be 6%?

OpenStudy (unklerhaukus):

sure that is a fine choice,

OpenStudy (anonymous):

24(1+.06)x?

OpenStudy (unklerhaukus):

So what do you get with those parameters if you plug them into \[f(x)=A(1+r)^x\]

OpenStudy (anonymous):

24(1+.06)^x

OpenStudy (unklerhaukus):

right , can you do the table on you own? i have to go to slepp

OpenStudy (anonymous):

sure thank you!

OpenStudy (unklerhaukus):

ok well, i hope i have helped you to understand some thing about exponential growth , good night

OpenStudy (anonymous):

you did! good night!